What is the Hardy-Winberge theorem and what is it’s importance to evolutionary

biology? Be detailed.
Answer #1

Allele and genotype frequencies in a population will remain constant from generation to generation in the absence of other evolutionary influences.

The biology defines evolution as being the sum total of the genetically inherited changes in the individuals who are the members of a population's gene pool. It is clear that the effects of evolution are felt by individuals, but it is the population as a whole that actually evolves. Evolution is simply a change in frequencies of alleles in the gene pool of a population.

For example, let us assume that there is a character that is determined by the inheritance of a gene with two alleles--A and B. If the parent generation has 92% A and 8% B and their offspring collectively have 90% A and 10% B, evolution has occurred between the generations. The entire population's gene pool has developed in the direction of a higher frequency of the B allele--it was not just those individuals who inherited the B allele who evolved.

This phenomena of evolution was given by Godfrey Hardy, an English mathematician, and Wilhelm Weinberg ia German physician in 20th Century with the help of mathematical modeling based on probability, they concluded in 1908 that gene pool frequencies are inherently stable but that evolution should be expected in all populations virtually all of the time. Both duo have concluded this apparent paradox by analyzing the net effects of potential evolutionary mechanisms.

Hardy, Weinberg, and the population geneticists who followed them came to understand that evolution will not occur in a population if seven conditions are met:

1. mutation is not occurring
2.   natural selection is not occurring
3.   the population is infinitely large
4.   all members of the population breed
5.   all mating is totally random
6.   everyone produces the same number of offspring
7.   there is no migration in or out of the population

These conditions are the absence of the things that can cause evolution.  In other words, if no mechanisms of evolution are acting on a population, evolution will not occur--the gene pool frequencies will remain unchanged. However, since it is highly unlikely that any of these seven conditions, let alone all of them, will happen in the real world, evolution is the inevitable result.

A duo went on to develop a simple equation that can be helpful to identify the probable genotype frequencies in a population and to ascertain their changes from one generation to another. This has become known as the Hardy-Weinberg equilibrium equation.  

In this equation (p² + 2pq + q² = 1), p is defined as the frequency of the dominant allele and q as the frequency of the recessive allele for a character controlled by a pair of alleles (A and a). In other words, p equals all of the alleles in individuals who are homozygous dominant (AA) and half of the alleles in people who are heterozygous (Aa) for this trait in a population. In mathematical terms, this is

p = AA + ½Aa

Likewise, q equals all of the alleles in individuals who are homozygous recessive (aa) and the other half of the alleles in people who are heterozygous (Aa).

q = aa + ½Aa

Because there are only two alleles in this case, the frequency of one plus the frequency of the other must equal 100%, which is to say

p + q = 1

Since this is logically true, then the following must also be correct:

p = 1 - q

There were only a few short steps from this knowledge for Hardy and Weinberg to realize that the chances of all possible combinations of alleles occurring randomly is

(p + q)² = 1

or more simply

p² + 2pq + q² = 1

In this equation, p² is the predicted frequency of homozygous dominant (AA) people in a population, 2pq is the predicted frequency of heterozygous (Aa) people, and q² is the predicted frequency of homozygous recessive (aa) ones.

From observations of phenotypes, it is usually only possible to know the frequency of homozygous recessive people, or q² in the equation, since they will not have the dominant trait. Those who express the trait in their phenotype could be either homozygous dominant (p²) or heterozygous (2pq). The Hardy-Weinberg equation allows us to predict which ones they are. Since p = 1 - q and q is known, it is possible to calculate p as well. Knowing p and q, it is a simple matter to plug these values into the Hardy-Weinberg equation (p² + 2pq + q² = 1). This then provides the predicted frequencies of all three genotypes for the selected trait within the population.